# Demystifying Network Foundation Models

### v2

@@ 4.5 Causal Sensitivity Testing: protocol-relevant perturbations
Table 5: Cosine similarity between baseline and embeddings obtained after perturbing IP and TCP header features.
@@
Table 5 presents results of our protocol-relevant perturbations testing on the CAIDA dataset, measuring how modifications to specific protocol features affect embedding stability. We selected CAIDA because it contains diverse, unfiltered Internet backbone traffic that represents real-world deployment, offering a more realistic evaluation of model robustness than controlled laboratory datasets.
Perturbation methodology accounts for model differences. Since NFMs employ substantially different tokenization techniques, we applied perturbations to semantically meaningful packet features (e.g., TCP Window Size, IP TTL) rather than model-specific tokens. Consequently, the same packet feature modification affects different percentages of tokens across models as shown in the “% tok” columns of Table 5. This approach ensures fair comparison by focusing on model-independent protocol semantics rather than model-specific implementations.
Models generally maintain stability under header perturbations. For most model-feature pairs (netFound, NetMamba, some features of YaTC), the cosine similarity between embeddings before and after header perturbation exceeds the average internal similarity of the baseline dataset. This indicates that single feature modifications do not significantly influence model representations from an embedding geometry perspective, indirectly showing architectural stability for header processing.
Selective feature sensitivity reveals architectural priorities. YaTC demonstrates selective sensitivity to transport layer features, particularly SEQ/ACK (0.61) and TCP Window Size (0.67), despite these modifications affecting minimal portions of the input (5% and 2.5% of tokens, respectively). This targeted sensitivity reveals architectural attention to specific protocol control signals, a pattern not evident in other models, which maintain near-baseline similarity across all header modifications.
Architectural differences in protocol processing. ET-BERT shows no sensitivity to header modifications as expected from its architecture, which deliberately ignores headers. netFound demonstrates remarkable stability (0.98-0.99) across all header modifications despite processing these fields, suggesting its representations prioritize higher-level patterns. These distinct responses directly reflect fundamental architectural differences in how models process protocol information.
Payload dependency across NFMs. Unlike header features, payload perturbations introduce significant representation shifts across almost all models (YaTC: 0.18, ET-BERT: 0.48, NetMamba: 0.62). These results reveal a critical dependency on payload features for decision-making, despite payload data often being encrypted or omitted in production environments. This finding highlights a potential robustness concern that wouldn’t be captured in traditional downstream evaluations.
@@ 4.6 Causal Sensitivity Testing: exogenous network context
We generated five sets of network traffic using a network emulator, NetReplica [15], each consisting of $\approx 100$ network flows with specific choices of Congestion Control, AQM, and used cross-traffic patterns. Please refer to Appendix J for additional details on NetReplica and generated datasets.
@@
Table 6 presents average cos similarity values for a diverse set of public datasets used in this work ("Average" row) and for the stability baseline dataset (a certain dataset with fixed parameters for all choices, selected as a baseline). Assuming that average and stability baseline represent lower and higher borders for possible similarity values for the particular model, for each different high-level context change (shown in different rows of the table as congestion control algorithm, AQM, and crosstraffic, or all for all changes), we present a $\Delta cos$ value which is calculated as the difference between the average cos similarity of the test dataset and the stability baseline. $F_{1}$ demonstrates the performance of a linear probe trained to separate between the stability baseline and the test dataset.
All models are able to group the baseline traffic. Table 6 (‘Average’ and ‘Stability baseline’ lines) shows that all models demonstrate much higher average coscos similarity inside the stability baseline compared to the average similarity in diverse datasets, which represents their ability to reliably group the similar traffic. This property is important for anomaly detection tasks, where previously unseen anomalous traffic should be separated from the normal traffic.
High-level context insignificantly influences the coscos similarity of embeddings. Given that average coscos similarity of diverse datasets and stability baseline are different for different models, we normalized $\Delta cos$ in Table 6 by the difference between the average coscos similarity of the stability baseline and average coscos similarity across diverse datasets, effectively measuring changes between the “lowest” and the “highest” measured coscos values. We notice that for all models, the $\Delta cos$ of synthetic datasets is pretty small (therefore, the resulting coscos is very close to the stability baseline). For almost all models and datasets (except NetMamba-AQM, NetMamba-All, and netFound-All), the absolute value of $\Delta cos$ is lower than 15%, showing models do not explicitly disentangle the high-level context from the baseline. In some cases, the $\Delta cos$ is even positive (therefore, grouping embeddings even closer), showing that the model considers the traffic with introduced changes closer to the stability baseline than the stability baseline itself, and therefore does not notice the high-level context change.
@@
Models easier separate AQM than Congestion Control, crosstraffic, or even all changes together. For almost all models, the linear probing $F_{1}$ score is significantly higher for the active queue management task compared to congestion control and crosstraffic. This shows that AQM influence is easier captured from the raw network traffic, and even sometimes easier than when all changes are introduced together. For all tasks, netFound demonstrates the highest $F_{1}$ score, being capable of capturing the high-level context changes. We also conducted additional measurements, including exploration of linear combinations of embeddings, in Appendix H.
@@ 5 Conclusion
In this paper, we introduce an initial approach to a principled evaluation of NFMs through intrinsic representation analysis and conduct a series of experiments on four SOTA NFMs over five diverse datasets for embedding geometry analysis, metric alignment assessment, and causal sensitivity testing. Our work reveals a number of critical insights, such as (1) all analyzed NFMs exhibit significant anisotropy that directly impacts downstream performance; (2) each model’s architectural design fundamentally determines which network metrics can be reliably extracted from its representations; and (3) all models demonstrate concerning sensitivity to payload information despite its frequent encryption in production environments. These findings highlight fundamental limitations in current NFM architectures and motivate future research on relationships between architectural choices, representation properties, and models’ performance.
@@ Appendix A Data representation for machine learning in networks
Network traffic (data exchanged between devices) is one of the main data sources for machine learning in networking. One of the unique characteristics of network traffic is that the same data is often represented in multiple ways simultaneously, such as raw packet captures that contain transmitted headers and data (PCAPs), flow records and statistics summarizing the single flow (finished communication session) between two endpoints (e.g., Zeek entries [39], IPFIX [4], NetFlow [14], CICFlowMeter [30], NTLFlowLyzer [45], etc.), and high-level extracted features (e.g., TCPInfo).
Being the most common representation, raw PCAPs contain all packets exchanged between the endpoints, including headers and payload. However, due to the nature of network communication and privacy concerns, the payload is often encrypted (e.g., TLS) and not available for analysis. Moreover, the scale of network traffic is so large that it is often impossible to analyze everything, and only a small subset of network flow is analyzed (e.g., the first several packets). All of this also limits labeled datasets availability, making the usage of unsupervised pretraining techniques beneficial.
@@ Appendix B Network Foundation Models and Datasets
ET-BERT [32] is a BERT-based network foundation model. ET-BERT uses only (encrypted) payload data of the first five packets of network traffic flow, ignoring packet headers, and is trained using masked token prediction and packet order prediction tasks. ET-BERT produces 768-dimensional embeddings.
netFound [23] is a transformer-based network foundation model that uses packet headers, payload, and additional flow information (such as packet inter arrival time or total bytes transferred), extracting packet burst representations (up to 60 packets), potentially from different parts of the traffic flow. In addition to masked token prediction, netFound uses multiple additional tasks, such as metadata prediction or packet order prediction, to pretrain the model. The used checkpoint of netFound (large) produces 1024-dimensional embeddings.
YaTC [59] and NetMamba [53] are masked autoencoder-based network foundation models (based on Transformer and Mamba architectures), which both transform existing network traffic data into a fixed-size image-like representation. Both models use the first five packets of network traffic flow, taking into account both headers and payload data. Models produce 192 and 256-dimensional embeddings, respectively.
@@
  • CIC-IDS2017 [46] is a benchmark intrusion detection dataset with 2.8M flows including both benign traffic and common attack patterns (brute force, DDoS, port scans, Heartbleed).
  • CIC-APT-IIoT24 [22] contains 3.7M flows representing 25 distinct advanced persistent threat techniques alongside benign traffic, testing model sensitivity to subtle attack patterns.
Real-World Traffic Datasets. These datasets were collected from production networks without artificial manipulation:
  • CAIDA [1] contains 1.5 million anonymized packet header flows from an Internet backbone link between Los Angeles and San Jose (collected in 2023).
@@ Appendix D Comparison of BatchNorm layer, decorrelation, and whitening
To further explore the influence of various normalization techniques, we implemented three different techniques (batch normalization, decorrelation, and whitening) and applied them to all models for all datasets containing downstream task labels. We trained a single linear layer for 30 epochs on the produced embeddings. Results are presented in Table 8. We notice that the decorrelation technique consistently provides the most valuable increase of $F_{1}$ test score among all measured, but also that the results might highly depend on the model and dataset combination.
@@ Appendix E CICFlowMeter feature correlation
Figure 2 provides an overview of the CKA similarity index for each feature produced by CICFlowMeter. The CKA values are averaged over all datasets. We noticed that ET-BERT demonstrates the lack of correlation between (payload-only) embeddings and the flow features, while YaTC and netFound clearly correlate with various features, some of which might be explicitly connected to their training algorithm and loss functions.
@@ Appendix F Manifold dimensionality
Models’ manifold ID is stable between datasets and similar between models Most of the models express similar manifold IDs between different datasets (except CIC-APT-IIoT24), indicating that the datasets contain approximately a similar amount of information extractable by the models, including both controlled (collected in the artificially established environment) and real-world datasets, and all models (except ET-BERT) are able to extract approximately the same amount of information from these datasets.
ET-BERT manifold ID represents the architectural choice. ET-BERT demonstrates much higher manifold ID values compared to all other models, which could be explained by the architectural choice of selecting only the payload for the embeddings computation. As the payload is often encrypted and therefore theoretically contains a uniformly distributed set of all possible input tokens, the model needs a much higher number of intrinsic dimensions to fully express the data.
@@ Appendix G Zero shot clusterization
We calculated the Silhouette score on the Crossmarket and CIC-IDS-2017 datasets for all models and presented the results in Table 10.
Most of the embeddings show a negative Silhouette score. We observe that for all datasets and models except ET-BERT and CIC-IDS-2017, the resulting score was negative. This indicates that the models’ extracted embeddings are grouped according to different patterns and similarities compared to those provided by the datasets’ authors. This demonstrates that embeddings extracted from the frozen pretrained models might not provide enough information for the downstream tasks, and additional unfreezing and training of the models might be required to improve their downstream-related extraction patterns.
@@
Inspired by Ethayarajh et al. [18], we designed an experiment to evaluate the linear transformation of the embedding space in networking foundation models. We used the same networking dataset as in subsection 4.6, where we use network traffic with a certain Congestion Control algorithm, AQM, and crosstraffic as a baseline $E_{Base}$, and the similar traffic with a different Congestion Control algorithm, AQM, and crosstraffic as $E_{CC}$, $E_{AQM}$, and $E_{Cross}$ respectively. We also define $E_{All}$ as a similar network traffic with all the different Congestion Control algorithms, AQMs, and crosstraffic from the baseline.
To investigate the linear transformation properties, we define the vector $E_{R}^{\prime}=E_{CC}+E_{AQM}+E_{Cross}$. As each of $E_{CC}$, $E_{AQM}$, and $E_{Cross}$ contain a single high-level context change and two unchanged contexts, to bring the resulting vector $E_{R}^{\prime}$ to $E_{All}$, we subtract $2*E_{Base}$ from $E_{R}^{\prime}$. Resulting vector $E_{R}$ is then defined as $E_{R}=E_{CC}+E_{AQM}+E_{Cross}-2*E_{Base}$ and should be close to $E_{All}$ if the linear transformation property holds.
@@
Table 11 shows the coscos similarity and $L_{1}$ distance from the centroid of the resulting embedding $E_{R}$ to the centroids of embeddings $E_{Base}$, $E_{CC}$, $E_{AQM}$, $E_{Cross}$, and $E_{All}$. Higher coscos similarity and lower $L_{1}$ distance indicate the closest embeddings to the resulting embedding $E_{R}$.
Neither of models seem to uphold the linear transformation property on both metrics. Among the four models, neither of models resulted in having the coscos similarity between $E_{R}$ and $E_{All}$ being the highest, with $E_{R}$ often aligning more closely with $E_{AQM}$ or $E_{Cross}$. At the same time, despite not having the highest coscos similarity, $E_{R}$ produced by netFound has the lowest $L_{1}$ distance to $E_{All}$, showing that the model is able (to some extent) to express the required linear transformation. Notably, both YaTC and netFound models have the lowest coscos similarity and the highest $L_{1}$ distance between $E_{R}$ and $E_{Base}$, which is an expected result as the resulting embedding $E_{R}$ should contain the most number of changes compared to the base embedding $E_{Base}$.
@@
NetReplica [15] is a programmable network emulator that offers realism, i.e., captures complex protocol dynamics and application behavior in generated data, and controllability, i.e., over tunable knobs to control the network conditions for data generation. It is originally designed to address the prevalent domain adaptation problem in networking, where models trained in (skewed) controlled settings fail to generalize in challenging production settings.
@@
System architecture. NetReplica decomposes a complex network environment into one or more bottleneck links (i.e., links that experience congestion) and facilitates explicitly specifying each of the links’ critical attributes to systematically create the type of network conditions. The system is built around three core abstractions: Link(type, node1, node2) for defining network paths, Bottleneck(type, link) for configuring constrained paths, and CrossTraffic for managing background traffic patterns. The system supports multiple backends, including LibreQoS, tc, and mahimahi for implementing traffic control policies.
@@
Specifically, it exposes two categories of control knobs that allow researchers to precisely specify network conditions for a bottleneck link: (1) Static attributes provide levers for specifying fundamental network conditions, including link capacity (e.g., 10 Mbps vs. 100 Mbps), maximum queue length (affecting bufferbloat), traffic shaping policy (e.g., token bucket or leaky bucket), and active queue management policy (e.g., Random Early Detection, Controlled Delay). (2) Dynamic attributes account for the behavior of the cross-traffic that traverses bottleneck links and leverages endogenously generated packet traces collected in situ from real production networks. These traces are preprocessed into different cross-traffic profiles (CTPs), each characterized by metrics measuring traffic intensity (average throughput), burstiness (temporal variation), and heterogeneity (traffic composition).
@@
Data-generation setup. We use three servers (A, B, C) with 2.2 GHz Intel Xeon processors, 192 GB RAM, and dual Intel-X710 10 Gbps NICs—one connected to the Internet via campus network, the other to a Cisco SX550X 10 Gbps switch. We configure the static and dynamic attributes for the bottleneck link that connects Server A and C (via Server B). We deploy video streaming (Puffer) and speedtest (NDT) servers on Server C with clients inside Docker containers on Server A. We use tcpreplay for replicating cross-traffic that offers precise timing control to preserve burst characteristics and temporal patterns during replay. The LibreQoS implementation at Server B uses an XDP-based bridge that provides efficient packet processing with minimal overhead.
@@
NetReplica-generated dataset. We use the combination of control applications at endpoints (i.e., Server A and C), static and dynamic attributes for the bottleneck link (LibreQoS at Server B) to synthesize different networking contexts. Each contextual subset contains $\approx 100$ network flows collected from the environment with fixed choices and small introduced variability in insignificant network metrics (e.g., minimal packet arrival time variation). In particular, this dataset contains the following parts:
@@
The cross-traffic profiles #29 and #36 were specifically selected to represent different combinations of intensity and burstiness. Profile #36 exhibits moderate intensity with low burstiness, while profile #29 features higher intensity with moderate burstiness, creating distinctly different network conditions. Each profile maintains consistent host counts and traffic composition to isolate the effects of temporal traffic patterns.

### v3

@@ 4.5 Causal Sensitivity Testing: protocol-relevant perturbations
Table 5: Cosine similarity between the baseline and embeddings after perturbing IP and TCP features.
@@
Table 5 presents results of our protocol-relevant perturbations testing on the CAIDA dataset, measuring how modifications to specific protocol features affect embedding stability. We selected the CAIDA dataset because it contains diverse, unfiltered, real-world Internet backbone traffic and allows therefore for a more realistic evaluation of model robustness than using controlled laboratory datasets.
Perturbation methodology accounts for model differences. Since the considered NFMs employ substantially different tokenization techniques, we applied perturbations to semantically meaningful packet features (e.g., TCP Window Size, IP TTL) rather than model-specific tokens. Consequently, the same packet feature modification affects different percentages of tokens across models as shown in the “% tok” columns of Table 5. This approach ensures fair comparison by focusing on model-independent protocol semantics rather than model-specific implementations.
Models generally maintain stability under header perturbations. For most model-feature pairs (netFound, NetMamba, some features of YaTC), the cosine similarity between embeddings before and after header perturbation exceeds the average internal similarity of the baseline dataset. These findings indicate that single feature modifications do not significantly influence model representations from an embedding geometry perspective, indirectly showing architectural stability for header processing.
Selective feature sensitivity reveals architectural priorities. YaTC demonstrates selective sensitivity to modifying transport layer features, particularly SEQ/ACK (0.61) and TCP Window Size (0.67), even though these modifications affect only small portions of the input (5% and 2.5% of tokens, respectively). This targeted sensitivity suggests architectural attention to specific protocol control signals, a pattern that is not evident in the other models, all of which maintain near-baseline similarity across all header modifications.
Architectural differences in protocol processing. ET-BERT shows no sensitivity to header modifications. This result is not surprising because the model’s architecture deliberately ignores headers. netFound demonstrates remarkable stability (0.98-0.99) across all header modifications despite processing these fields, suggesting its representations prioritize higher-level patterns. These distinct responses directly reflect fundamental differences in how models process protocol information.
Payload dependency across NFMs. Unlike header features, payload perturbations introduce significant representation shifts across almost all models (YaTC: 0.18, ET-BERT: 0.48, NetMamba: 0.62). These results reveal a critical dependency on payload features for decision-making, even though this data is often encrypted or omitted in production environments. These findings highlight a potential model robustness concern that would go unnoticed in traditional downstream evaluations.
@@ 4.6 Causal Sensitivity Testing: exogenous network context
We use the network emulator NetReplica [15] to generated five synthetic datasets, each consisting of $\approx 100$ network flows and obtained using specific choices of congestion control algorithms, AQM policies, and cross-traffic patterns. Additional details about NetReplica and the generated datasets can be found in Appendix J.
@@
Table 6 presents average cos similarity values for a diverse set of public datasets used in this work ("Average" row) and for the stability baseline dataset (i.e., a fixed dataset with given parameters for all choices, selected as a baseline). Assuming that the average and stability baselines represent lower and higher bounds respectively for the possible similarity values for each model, we list for each different high-level context change (see rows of the table labeled "Congestion Control", "AQM", "Crosstraffic", and "All" for all changes) the $\Delta cos$ value which is calculated as the difference between the average cos similarity of the test dataset and the stability baseline. The reported $F_{1}$ scores denote the performance of a linear probe trained to separate between the stability baseline and the test dataset.
All models are able to group the baseline traffic. Table 6 (‘Average’ and ‘Stability baseline’ lines) shows that all models demonstrate much higher average coscos similarity (but staying below the upper bound provided by the stability baseline) compared to the average similarity in diverse datasets. This property suggest that the models are able to reliably group the similar traffic and is important, for example, for anomaly detection tasks, where previously unseen anomalous traffic should be separated from the normal traffic.
High-level context insignificantly influences the coscos similarity of embeddings. Given that average coscos similarity of diverse datasets and stability baseline are different for different models, we normalized $\Delta cos$ in Table 6 by the difference between the average coscos similarity of the stability baseline and average coscos similarity across diverse datasets, effectively measuring changes between the “lowest” and the “highest” measured coscos values. We notice that for all models, the $\Delta cos$ of the synthetic datasets is relatively small, implying that the resulting coscos is very close to the stability baseline. Except for NetMamba-AQM, NetMamba-All, and netFound-All, for all other model-dataset pairs, the absolute value of $\Delta cos$ is lower than 15%, an indication that the models do not explicitly disentangle the high-level context from the baseline. In some cases, the $\Delta cos$ is even positive (i.e., embeddings are grouped even closer), demonstrating that the model considers the traffic with the introduced changes to be more similar to the stability baseline than the stability baseline itself and does therefore not notice the high-level context change.
@@
Models are better at distinguishing different AQMs than different Congestion Control algorithms, crosstraffic patterns, or even all changes together. For almost all models, the linear probing $F_{1}$ score is significantly higher for the AQM row in Table 6 compared to the Congestion Control and Crosstraffic rows. This observation suggests that the influence of AQM policies is easier captured from the raw network traffic, and sometimes even easier than when all changes are introduced together. Across all four rows, netFound has the highest $F_{1}$ scores, an indication that it is capable of capturing the high-level context changes. We also conducted additional experiments, including examining linear combinations of embeddings, and report the results in Appendix H.
@@ 5 Conclusion
In this paper, we introduce an initial approach to a principled evaluation of NFMs through intrinsic representation analysis and conduct a series of experiments on four SOTA NFMs over five diverse datasets for embedding geometry analysis, metric alignment assessment, and causal sensitivity testing. Our work reveals a number of critical insights, such as (1) all analyzed NFMs exhibit significant anisotropy that directly impacts downstream performance; (2) each model’s architectural design fundamentally determines which network metrics can be reliably extracted from its representations; and (3) all models demonstrate concerning sensitivity to payload information despite its frequent encryption (or omission) in production environments. In particular, our proposed framework provides the network community with systematic tools for comparing different NFM architectures objectively, identifying model limitations before deployment, guiding architectural improvements based on intrinsic properties, and understanding failure modes that downstream metrics miss. At the same time, our findings highlight fundamental limitations in current NFM architectures and motivate future research on relationships between architectural choices, representation properties, and models’ performance.
@@ Appendix A Data representation for machine learning in networks
Network traffic (data exchanged between devices) is one of the main data sources for machine learning in networking. One of the unique characteristics of network traffic is that the same data is often represented in multiple ways simultaneously, such as raw packet captures that contain transmitted headers and payload (PCAPs), flow records and statistics summarizing each individual flow (finished communication session) between two endpoints (e.g., Zeek entries [39], IPFIX [4], NetFlow [14], CICFlowMeter [30], NTLFlowLyzer [45], etc.), and high-level extracted features (e.g., TCPInfo).
Being the most common representation, raw PCAPs contain all packets exchanged between the endpoints, including headers and payload. However, due to the nature of network communication and privacy concerns, the payload is often encrypted (e.g., TLS) and not available for analysis. Moreover, the volume of network traffic is often so large that it is difficult to analyze everything. In such cases, a practical solution is to analyze only a small portion of each network flow (e.g., the first few packets). These and other challenges also limit the availability of labeled datasets, making the use of self-supervised pretraining techniques especially attractive.
@@ Appendix B Network Foundation Models and Datasets
ET-BERT [32] is a BERT-based network foundation model. ET-BERT uses only (encrypted) payload data of the first five packets of network traffic flows, ignoring packet headers, and is trained using masked token prediction and packet order prediction tasks. ET-BERT produces 768-dimensional embeddings.
netFound [23] is a transformer-based network foundation model that uses packet headers, payload, and additional flow information (such as packet interarrival times or total bytes transferred), extracting packet burst representations (up to 60 packets), potentially from different parts of the traffic flow. In addition to masked token prediction, netFound uses multiple additional tasks, such as metadata prediction or packet order prediction, to pretrain the model. The used checkpoint of netFound (large) produces 1024-dimensional embeddings.
YaTC [59] and NetMamba [53] are masked autoencoder-based network foundation models (based on Transformer and Mamba architectures), which both transform existing network traffic data into a fixed-size image-like representation. Both models use the first five packets of each network traffic flow, taking into account both headers and payload data. The two models produce 192 and 256-dimensional embeddings, respectively.
@@
  • CIC-IDS2017 [46] is a benchmark intrusion detection dataset with 2.8M flows including both benign traffic and common attack patterns (e.g., brute force, DDoS, port scans, Heartbleed).
  • CIC-APT-IIoT24 [22] contains 3.7M flows representing 25 distinct advanced persistent threat techniques alongside benign traffic and is valuable for testing model sensitivity to subtle attack patterns.
Real-World Traffic Datasets. These datasets were collected from actual production networks without artificial manipulation:
  • CAIDA [1] contains 1.5 million anonymized packet-level flows from an Internet backbone link between Los Angeles and San Jose (collected in 2023).
@@ Appendix C Anisotropy of the embedding space
Table 8: MCC distribution statistics across all datasets.
Model Range Std Skew Kurtosis Gini
YaTC 0.2105 0.0163 10.7569 129.9826 0.7522
ET-BERT 0.0157 0.0017 3.3156 16.2447 0.6547
netFound 0.0119 0.0012 3.5168 19.9052 0.6504
NetMamba 0.0250 0.0049 2.1100 4.4845 0.6388
Table 8 lists MCC distribution-related statistics across all datasets, providing additional insight into the variability of the MCC metric.
@@ Appendix D Comparison of BatchNorm layer, decorrelation, and whitening
To further explore the influence of various normalization techniques, we implemented three different techniques (batch normalization, decorrelation, and whitening) and applied them to all models for all datasets containing downstream task labels. We trained a single linear layer for 30 epochs on the produced embeddings. The results are presented in Table 9 and show that the decorrelation technique consistently provides the largest increase in the $F_{1}$ test scores among all implemented techniques. At the same time, we also notice that the results can be expected to depend on the considered model-dataset combination.
@@ Appendix E CICFlowMeter feature correlation
Figure 3 provides an overview of the CKA similarity index for each feature produced by CICFlowMeter. The CKA values are averaged over all datasets. We notice that ET-BERT shows a lack of correlation between (payload-only) embeddings and the flow features, while YaTC and netFound clearly correlate with various features, some of which might be explicitly connected to their training algorithm and loss functions.
@@ Appendix F Manifold dimensionality
Models’ manifold ID is stable between datasets and similar between models. Most of the models have similar manifold IDs between different datasets (except CIC-APT-IIoT24), indicating that the datasets contain approximately similar amounts of information extractable by the models, irrespective of whether the dataset is synthetic (collected in a controlled environment) or realistic (collected in a real-world production network). Also, all models (except ET-BERT) are able to extract approximately the same amount of information from these datasets.
ET-BERT manifold ID represents an architectural choice. ET-BERT demonstrates much higher manifold ID values compared to all other models. This observation can be explained by the model’s architectural choice of selecting only the payload for the embeddings computation. Since the payload is often encrypted and contains therefore in theory a uniformly distributed set of all possible input tokens, the model needs a much higher number of intrinsic dimensions to fully express the data.
@@ Appendix G Zero shot clusterization
We calculated the Silhouette score on the Crossmarket and CIC-IDS-2017 datasets for all models and present the results in Table 11.
@@
Most of the embeddings have a negative Silhouette score. We observe that for both datasets and all models (except ET-BERT and CIC-IDS-2017), the resulting score is negative. This finding suggests that the models’ extracted embeddings are grouped according to different patterns and similarities compared to those provided by the datasets’ authors. It also shows that embeddings extracted from the frozen pretrained models might not provide enough information for the downstream tasks and that additional unfreezing and training of the models might be required to improve their downstream-related extraction patterns.
@@
Inspired by Ethayarajh et al. [18], we designed an experiment to evaluate the linear transformation of the embedding space in networking foundation models. We used the same networking datasets as in subsection 4.6, where we use network traffic with a certain Congestion Control algorithm, AQM policy, and crosstraffic patterns for the baseline $E_{Base}$, and similar kind of traffic but with a different Congestion Control algorithm, AQM policy, and crosstraffic pattern for $E_{CC}$, $E_{AQM}$, and $E_{Cross}$ respectively. We also define $E_{All}$ for the network traffic with all the different Congestion Control algorithms, AQM policies, and crosstraffic patterns from the baseline.
To investigate the linear transformation properties, we define the vector $E_{R}^{\prime}=E_{CC}+E_{AQM}+E_{Cross}$. As each of $E_{CC}$, $E_{AQM}$, and $E_{Cross}$ contains a single high-level context change and two unchanged contexts, to bring the resulting vector $E_{R}^{\prime}$ to $E_{All}$, we subtract $2*E_{Base}$ from $E_{R}^{\prime}$. The resulting vector $E_{R}$ is then defined as $E_{R}=E_{CC}+E_{AQM}E_{Cross}-2*E_{Base}$ and should be close to $E_{All}$ if the linear transformation property holds.
Table 12 shows the coscos similarity and $L_{1}$ distance from the centroid of the resulting embedding $E_{R}$ to the centroids of the embeddings $E_{Base}$, $E_{CC}$, $E_{AQM}$, $E_{Cross}$, and $E_{All}$. Higher coscos similarity and lower $L_{1}$ distance indicate the closest embeddings to the resulting embedding $E_{R}$.
Neither of models seem to uphold the linear transformation property in both metrics. Among the four models, neither of the models managed to demonstrate the desired result of coscos similarity being the highest between $E_{R}$ (resulting embedding after linear transformation) and $E_{All}$ (target embedding), instead often aligning $E_{R}$ more closely with $E_{AQM}$ or $E_{Cross}$. At the same time, $E_{R}$ produced by netFound has the lowest $L_{1}$ distance to $E_{All}$, showing that the model is able to express (to some extent) the required linear transformation properties and contains correct internal mapping of network context in the embedding space.
In addition, both the YaTC and netFound models have the lowest coscos similarity and the highest $L_{1}$ distance between $E_{R}$ and $E_{Base}$ (original embedding before any changes or transformations). This result highlights that both YaTC and netFound can notice high-level network context changes introduced by Congestion Control, AQM, and crosstraffic variations, distancing the resulting embedding from the original traffic even if this distance does not uphold linear transformation properties (in case of YaTC).
@@
NetReplica [15] is a programmable network emulator that offers realism (i.e., captures complex protocol dynamics and application behavior in the generated data), and controllability (i.e., provides tunable knobs to control the network conditions for data generation). The emulator was originally designed to address the prevalent domain adaptation problem in networking, where models trained in (skewed) controlled settings fail to generalize in challenging production settings.
@@
System architecture. NetReplica decomposes a complex network environment into one or more bottleneck links (i.e., links that experience congestion) and facilitates explicitly specifying each of the links’ critical attributes to systematically create different types of network conditions. The system is built around three core abstractions: Link(type, node1, node2) for defining network paths, Bottleneck(type, link) for configuring constrained paths, and CrossTraffic for managing background traffic patterns. The system supports multiple backends, including LibreQoS, tc, and mahimahi for implementing traffic control policies.
@@
Specifically, NetReplica exposes two categories of control knobs that allow researchers to precisely specify network conditions for a bottleneck link: (1) Static attributes provide levers for specifying fundamental network conditions, including link capacity (e.g., 10 Mbps vs. 100 Mbps), maximum queue length (affecting bufferbloat), traffic shaping policy (e.g., token bucket or leaky bucket), and active queue management policy (e.g., Random Early Detection, Controlled Delay). (2) Dynamic attributes account for the behavior of the cross-traffic that traverses bottleneck links and leverages endogenously generated packet traces collected in situ from real production networks. These traces are preprocessed into different cross-traffic profiles (CTPs), each characterized by metrics measuring traffic intensity (average throughput), burstiness (temporal variation), and heterogeneity (traffic composition).
@@
Data-generation setup. We use three servers (A, B, C) with 2.2 GHz Intel Xeon processors, 192 GB RAM, and dual Intel-X710 10 Gbps NICs – one connected to the Internet via our campus network, the other to a Cisco SX550X 10 Gbps switch. We configure the static and dynamic attributes for the bottleneck link that connects Server A and C (via Server B). We deploy video streaming (Puffer) and speedtest (NDT) servers on Server C with clients inside Docker containers on Server A. We use tcpreplay for replicating cross-traffic that offers precise timing control to preserve burst characteristics and temporal patterns during replay. The LibreQoS implementation at Server B uses an XDP-based bridge that provides efficient packet processing with minimal overhead.
@@
NetReplica-generated dataset. We use a combination of control applications at endpoints (i.e., Server A and C) and static and dynamic attributes for the bottleneck link (LibreQoS at Server B) to synthesize different networking contexts. Each contextual subset contains $\approx 100$ network flows collected from the environment with fixed choices and small introduced variability in insignificant network metrics (e.g., minimal packet arrival time variation). In particular, this dataset contains the following subsets:
@@
The cross-traffic profiles #29 and #36 were purposefully selected to represent different combinations of intensity and burstiness. Profile #36 exhibits moderate intensity with low burstiness, while profile #29 features higher intensity with moderate burstiness, creating distinctly different network conditions. Each profile maintains consistent host counts and traffic composition to isolate the effects of temporal traffic patterns.